20 points Problem 4: Extended Euclidean Algorithm Using Extended Euclidean Algorithm compute the greatest common divisor and(c) a 4930 and b-1292 (d) a -7540 and b 2795
Answer #1

P-L a-270 , b= 21の 270-3x2-15-ナ 51, . aed (2.705.219) . 2x 51 21の一 2.2.x (270-21の)-, 2x270 の 3721acal (86の,605)=1 dl @ a-4の3o , = 12の2. god (4989, 12.02 54 = 44238-102 - 2 4-210532 の×298-2x1D54 군24x26 ヌ 27 5-1956),- 3x1の50 ニマ乂2マの5-10X1950. Bezouts a1o2